OpenStudy (anonymous):

What is an equation of the line that passes through (0, 8) and (4, 0)? y = -2x + 4 y = 2x + 4 y = -2x + 8 y = 2x + 8

OpenStudy (anonymous):

Find the slope first with (y2-y1)/(x2-x1)

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

Do you know how to find slope?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

We have point (x1,y1) which is (0,8) and point (x2,y2) which is (4,0) Plug the into the equation I posted above

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

(y2-y1)/(x2-x1) (0-8)/(4-0) = ?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

-8 ---- = -2 4

OpenStudy (anonymous):

so whqts the answer i will give you a medal

OpenStudy (anonymous):

Now we have to plug -2 (slope) into another equation. I will choose point (0,8). y-y1=m(x-x1); where y1 is 8, m is -2, and x1 is 0

OpenStudy (anonymous):

so its d?

OpenStudy (anonymous):

y-8=-2(x-0) can you solve this for y?

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

y-8=-2(x-0) y-8=-2x +8 +8 y=-2x+8